OpenStudy (anonymous):

Solve for x: 1 over 4 (2x − 14) = 4 15 15 over 2 9 11

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }\times \left( 2x-14 \right)=4\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }\times2x-\frac{ 1 }{ 4 }\times14=4\]

OpenStudy (anonymous):

\[\frac{ 2x }{ 4 }-\frac{ 14 }{ 4 }=4\]

OpenStudy (anonymous):

\[\frac{ 2x }{ 4 }=4+\frac{ 7 }{ 2 }\]

OpenStudy (anonymous):

\[x=4+7\]

OpenStudy (anonymous):

x=11